# EXGnet: a single-lead explainable-AI guided multiresolution network with train-only quantitative features for trustworthy ECG arrhythmia classification

### v1

@@ Abstract
Background: Deep learning has significantly advanced ECG arrhythmia classification, enabling high accuracy in detecting various cardiac conditions. The use of single-lead ECG systems is crucial for portable devices, as they offer convenience and accessibility for continuous monitoring in diverse settings. However, the interpretability and reliability of deep learning models in clinical applications poses challenges due to their black-box nature.
Methods: To address these challenges, we propose EXGnet, a single-lead, trustworthy ECG arrhythmia classification network that integrates multiresolution feature extraction with Explainable Artificial Intelligence (XAI) guidance and train only quantitative features.
Results: Trained on two public datasets, including Chapman and Ningbo, EXGnet demonstrates superior performance through key metrics such as Accuracy, F1-score, Sensitivity, and Specificity. The proposed method achieved average five fold accuracy of 98.762%, and 96.932% and average F1-score of 97.910%, and 95.527% on the Chapman and Ningbo datasets, respectively.
Conclusions: By employing XAI techniques, specifically Grad-CAM, the model provides visual insights into the relevant ECG segments it analyzes, thereby enhancing clinician trust in its predictions. While quantitative features further improve classification performance, they are not required during testing, making the model suitable for real-world applications. Overall, EXGnet not only achieves better classification accuracy but also addresses the critical need for interpretability in deep learning, facilitating broader adoption in portable ECG monitoring.
keywords:
single-lead ECG systems, ECG arrhythmia classification, trustworthy, Explainable Artificial Intelligence, XAI guided, Grad-CAM
@@ 1 Introduction
Cardiovascular disease (CVD) is the leading cause of death globally, affecting over half a billion people, with 20.5 million deaths in 2021 [1]. More than 80% of these fatalities result from heart attacks and strokes, with one-third occurring prematurely in individuals under 70 years old [2]. These CVDs can be detected and monitored via various diagnostic tools. One of the most effective is the electrocardiogram (ECG), a fast and simple test that measures heart rate, rhythm, and timing by recording natural electrical impulses in the heart’s various parts [3]. Recent advancements in ECG technology have enabled the developments of portable, cost-effective, and user-friendly devices for point-of-care screening and continuous monitoring. These technologies mostly utilize single-lead ECGs, simplifying the monitoring process of heart activity.
@@
Within the evolving landscape, two main categories of CVD detection technologies that have emerged are traditional methods and deep learning-based methods. Traditional methods comprise of expert features extraction stage and a machine learning algorithm prediction stage. Support vector machine (SVM) classifiers are widely utilized among machine learning methods [4]. Notably, Ye et.al.[5] proposed integrating SVM classifiers from ECG samples with wavelet coefficient and RR interval. Discrete wavelet transformation (DWT) was used in early research to extract the QRS wave’s morphological characteristics. Wavelet packet decomposition (WPD) was proposed as an extension to classical DWT because it can only decompose low-frequency components (approximations). This allows both high-frequency and low-frequency components to be used [6]. Chen et.al. [7] developed an SVM classifier utilizing RR intervals and ECG projection matrices to classify variable-length ECG signals. Shi et.al. [8] used XGBoost algorithms that work with handcraft features to classify ECG, but the model struggled with accurately predicting abnormal classes. Another study incorporated machine learning models including SVM, KNN, and ANN for classification and included parametric features—more precisely, the interval, amplitude, and length of ECG morphology—to improve performance [9]. Cyclic spectral analysis was employed in conjunction with a generalized autoregressive conditional heteroskedasticity model, followed by SVM-based classification, to enhance the analysis of ECG signals by Mihandoost and Amirani [10]. The main drawback of these traditional methods is that they require hand-crafted features that are dependent on proper ECG signal processing and human intervention. Low-quality or noisy signal fails to provide proper features effectively. Furthermore, experimenting with various feature reduction techniques is essential to identify the optimal set of features.
Table 1: Summary of recent studies on arrhythmia classification using deep learning algorithms.
Author Year Class Approach Database Accuracy Sensitivity Specificity
Hannun et.al. [11] 2019 AFib-AF, AVB, EAR, IVR, Junctional, Noisy, SR, SVT, VT, Trigeminy, Bigeminy CNN (Private) - 75.22 97.2
Petmezas et.al. [12] 2021 AFib, NSR, AF, J CNN, LSTM MIT-BIH 97.4 97.0 98.4
Guan et.al. [13] 2022 SB, SR, SI AFL, AFib ST, SVT HA-Resnet Chapman 88.16 79.08 98.16
Serhal et.al. [14] 2023 AFib, NSR EMD, CNN PTB-XL 98.80 - -
Hao et.al. [15] 2023 SR, SB, ST, AF, AFib, SVT G2-ResNeXt Chapman 98.38 97.34 99.59
Xie et.al. [16] 2024 AFib, non-AFib CWT-MB-ResNet Physionet - - -
Singhal et.al. [17] 2024 AFib, SR, VF GSMD-SRST Physionet, Mendeley 99.20 98.60 99.00
Gais et.al. [18] 2024 N, S, V, F, Q Bidirectional Transformer MIT-BIH 94.26 93.15 -
Choi et.al. [19] 2024 AFib, non-AFib LSTM, XGBoost PTB-XL, Chapman 93.00 95.35 89.56
Chen et.al. [20] 2024 Normal, AFib, Other, Noisy EMCNet Physionet 90.50 85.00 92.00
Goettling et.al. [21] 2024 AFib, non-AFib xECGArch PTB-XL, Georgia, CPSC 2018, Chapman 95.33 94.87 95.82
Ayano et.al. [22] 2024 SR, SB, ST, AF-AFib, SVT CNN-BiLSTM Chapman 95.81 93.79 98.94
@@
To overcome these limitations of traditional approaches, deep CNNs have recently gained significant attention from researchers due to their powerful ability to automatically extract and select features directly from input signals. Asgharzadeh et.al. [23] proposed calculating spectral entropy to generate a three-dimensional time-frequency transform for QRS complexes and employed bi-directional, two-dimensional PCA to filter out irrelevant features and enhance class separation by simultaneously applying PCA to the rows and columns of the TFT followed by a CNN for classification. Building on this, a novel hidden attention residual network was proposed to extract useful features from ECG signals by converting 2s segments into images [13]. Additionally, a lightweight fusing transformer model was introduced to address challenges related to large-scale parameters and performance degradation, further advancing the capabilities of ECG analysis [24]. Tao et.al. [25] incorporated an expert-knowledge attention network to recognize four tachyarrhythmias, highlighting the need for deep learning approaches to consider inter-channel shared information and time-sequence dependent signals. In another study [26], a lightweight custom CNN and transfer learning model with MobileNet-V2 architecture was applied to segmented ECG image data that was balanced by auxiliary classifier generative adversarial network (ACGAN). Chen et.al. [20] developed an ensemble multiscale CNN with two modules: LSTM network learns from 1-D signals and time-frequency spectrograms from denoised ECG signals on VGGNet. The CWT scalogram from a single-lead ECG signal was employed in a study by Xie et.al. [16]. In a different research, spectrograms of 1D-ECG signals were created using the superlet-transform-based (SLT) method as opposed to the CWT and ResNet-18 for classification [17]. Islam et.al. [27] proposed a attention, convolution, and transformer-based network, to effectively capture heartbeat morphological characteristics using peak annotation on ECG signals incorporating preprocessing algorithms like synthetic minority oversampling technique edited nearest neighbor (SMOTE-ENN), SMOTE-Tomek, and adaptive synthetic. On another approach, a novel inter-patient heartbeat classification model, G2-ResNeXt, was proposed to enhance the recognition of patient’s heart conditions by incorporating a two-fold grouping convolution (G2) into the original ResNeXt structure [15]. To tackle class imbalances in ECG data, a context-aware loss was introduced in [18], which dynamically adjusts weights according to class representation.
@@
Explainable artificial intelligence (XAI) methods are relatively new in the field of deep learning-based biomedical signal analysis, with only a few studies exploring various XAI techniques for ECG classification. Jo et al. [28] developed a CNN architecture with submodules to detect the presence of P waves and RR irregularities, using gradient class activation maps (Grad-CAM) to interpret the results. Bender et.al. [29] employed multiple layer-wise relevance propagation (LRP) rules and integrated gradients (IG) for interpreting different arrhythmias. In contrast, Singh & Sharma [30] performed a more comprehensive comparison of four XAI methods, including Grad-CAM, and Shapley Additive Explanations (SHAP). Goettling et.al. [21] compared 13 different XAI methods, and among those techniques, the deep taylor decomposition (DTD) explanation seemed more trusted. H. Le et.al. [31] used modified Grad-CAM for lead-wise explanation. They instantiate Spearman’s rank correlation to assess the quality of the XAI approach. In another work, Ayano et.al. [22] employed two model interpretability techniques namely, Grad-CAM++ and SHAP. Table 1 contains a brief overview of works along with their corresponding classification performance scores.
Most studies using deep CNNs, RNNs, or hybrid methods have solely trained on raw ECG signals or expert features. However, certain heart arrhythmias have correlations with specific morphological features, making it essential to consider both for accurate interpretation. To improve performance, some studies have integrated these quantitative features with deep learning networks. Besides, in many cases, the visualizations generated by XAI techniques may not align with verbal descriptions used in medical practice and may lack clarity in highlighting key ECG segments. We address these limitations by introducing a novel architecture, EXGnet, which provides a reliable, trustworthy, and interpretable deep learning framework for ECG analysis. Our major contributions are as follows:
@@
  • We introduce a novel deep learning architecture that efficiently extracts multiresolution features from ECG signal while minimizing computational complexity. This approach enables the model to learn both short-term and long-term features within a single network.
@@
  • We introduce XAI guidance through a Normalized Cross-Correlation (NCC)-based loss function to guide Class Activation Map (CAM) values. This novel approach ensures that our deep learning model focus on the specific regions of ECG signals that cardiologists typically examine for decision-making.
@@
  • We strategically integrate quantitative features solely during the training phase of our network, enabling it to leverage these features for robust learning. Notably, during testing, no quantitative features are needed, yet the network delivers improved performance compared to when these features were absent during training.
@@ 2.1.1 Datasets
To evaluate the performance of EXGnet, we conducted experiments using the publicly available database from the PhysioNet/Computing in Cardiology Challenge 2021 [32]. This database includes ECG recordings from various challenges and hospitals, totaling 45,512 recordings from the Chapman-Shaoxing [33] (Chapman dataset) and Ningbo [34] datasets, each sampled at 500 Hz for 10 seconds. Both of these datasets are jointly available online in a Physionet official link (https://physionet.org/content/ecg-arrhythmia/1.0.0/).
We focused on the following heart conditions: Sinus Rhythm (SR), Sinus Arrhythmia (SA), Sinus Tachycardia (ST), Sinus Bradycardia (SB), Atrial Fibrillation (AFib), Atrial Flutter (AF), Atrial Tachycardia (AT), and Supraventricular Tachycardia (SVT). These were selected for effective identification using the lead II signals. To enable multiclass classification, we ensured each ECG recording had a unique class, although some signals may belong to multiple categories. This resulted in 5,086 recordings from the Chapman dataset and 16,411 from the Ningbo dataset. Due to insufficient data, we combined AF with AFib and AT with SVT. Besides, research shows that AF and AFib signals are often quite similar in these datasets [35]. The class distribution is detailed in Table 2. Since there are only 15 samples available for the combined AT and SVT class in the Ningbo dataset, we decided to omit this class from our analysis.
@@ 2.1.2 Quantitative features extraction
In our work, we utilized quantitative features extracted from ECG signals using the NeuroKit 0.2 Python library [36]. This library offers a comprehensive set of functions that can be applied to ECG data to extract essential features, aiding in detailed signal analysis. The following 17 features have been extracted from the ECG signal: beats per minute (BPM), mean of normal-to-normal intervals (Mean NN), standard deviation of successive differences (SDSD), standard deviation of NN intervals (SDNN), and root mean square of successive differences (RMSSD). Additional features include the coefficient of variation of successive differences (CVSD), coefficient of variation of NN intervals (CVNN), median NN intervals (Median NN), median absolute deviation of NN intervals (MAD NN), and the median coefficient of variation of NN intervals (MCV NN). We also extracted the interquartile range of NN intervals (IQR NN), the standard deviation of the RMSSD (SDRMSSD), 20th and 80th percentiles of NN intervals (Prc20 NN, Prc80 NN), minimum NN intervals (Min NN), maximum NN intervals (Max NN), and the heart rate turbulence index (HTI).
@@ 2.1.3 Data preprocessing
Figure 2: Detailed block diagram of the proposed EXGnet

### v2

@@ Abstract
Deep learning has significantly propelled the performance of ECG arrhythmia classification, yet its clinical adoption remains hindered by challenges in interpretability and deployment on resource-constrained edge devices. To bridge this gap, we propose EXGnet, a novel and reliable ECG arrhythmia classification network tailored for single-lead signals, specifically designed to balance high accuracy, explainability, and edge compatibility. EXGnet integrates explainable artificial intelligence (XAI) supervision during training via a normalized cross-correlation based loss, directing the model’s attention to clinically relevant ECG regions, similar to a cardiologist’s focus. This supervision is driven by automatically generated ground truth, derived through an innovative heart rate variability-based approach, without the need for manual annotation. To enhance classification accuracy without compromising deployment simplicity, we incorporate quantitative ECG features during training. These enrich the model with multi-domain knowledge but are excluded during inference, keeping the model lightweight for edge deployment. Additionally, we introduce an innovative multiresolution block to efficiently capture both short- and long-term signal features while maintaining computational efficiency. Rigorous evaluation on the Chapman and Ningbo benchmark datasets validates the supremacy of EXGnet, which achieves average five-fold accuracies of 98.762% and 96.932%, and F1-scores of 97.910% and 95.527%, respectively. Comprehensive ablation studies and both quantitative and qualitative interpretability assessment confirm that the XAI guidance is pivotal, demonstrably enhancing the model’s focus and trustworthiness. Overall, EXGnet sets a new benchmark by combining high-performance arrhythmia classification with interpretability, paving the way for more trustworthy and accessible portable ECG based health monitoring systems.
Keywords Single-lead ECG systems  $\cdot$ ECG arrhythmia classification  $\cdot$ Trustworthy  $\cdot$ Explainable Artificial Intelligence  $\cdot$ XAI guidance
@@ 1 Introduction
Cardiovascular disease (CVD) is the leading cause of death globally, affecting over half a billion people, with 20.5 million deaths in 2021 [1]. More than 80% of these fatalities result from heart attacks and strokes, with one-third occurring prematurely in individuals under 70 years old [2]. These CVDs can be detected and monitored using various diagnostic tools. One of the most effective tools is the electrocardiogram (ECG), a fast and simple test that measures heart rate, rhythm, and timing by recording natural electrical impulses in the heart’s various parts [3]. Recent advances in deep learning and edge computing have led to cost-effective and user-friendly portable ECG devices, especially single-lead systems, which strongly support smartphone-based ECG analysis at home and empower individuals to monitor their heart health conveniently and continuously [4].
@@
Within the evolving landscape, two main categories of CVD detection technologies have emerged: traditional methods and deep learning-based methods. Traditional methods comprise an expert feature extraction stage and a machine learning algorithm prediction stage. Support Vector Machine (SVM) classifiers are widely utilized among machine learning methods [5]. Early research used Discrete Wavelet Transform (DWT) to extract QRS morphology, but its limitation to low-frequency decomposition led to Wavelet Packet Decomposition (WPD), which includes both high- and low-frequency components [6]. Recent work has demonstrated that combining Time Series Feature Extraction Library (TSFEL)-based feature extraction with balanced ensemble learning, specifically a weighted majority vote of Random Forest and SVM classifiers, can significantly improve arrhythmia classification accuracy on the MIT-BIH database [7]. The main drawback of these traditional methods is that they require hand-crafted features that are dependent on proper ECG signal processing and human intervention. A low-quality or noisy signal fails to provide proper features effectively. Furthermore, experimenting with various feature reduction techniques is essential to identify the optimal set of features.
Figure 1: Overview of ECG healthcare system where EXGnet can be integrated for reliable real-time heart arrhythmia detection.
@@
To overcome the limitations of traditional approaches, deep Convolutional Neural Networks (CNNs) have gained significant attention from researchers due to their powerful ability to automatically extract and select features directly from input signals. Asgharzadeh et al. [8] proposed calculating spectral entropy to generate a three-dimensional time-frequency transform for QRS complexes and employed bi-directional, two-dimensional Principal Component Analysis (PCA) to filter out irrelevant features and enhance class separation by simultaneously applying PCA to the rows and columns of the Short-Time Fourier Transform (STFT), followed by a CNN for classification. Based on this, a novel hidden attention residual network was proposed to extract useful features from ECG signals by converting 2-second segments into images [9]. Additionally, a lightweight fusing transformer model was introduced to address challenges related to large-scale parameters and performance degradation, further advancing the capabilities of ECG analysis [10]. Tao et al. [11] incorporated an expert-knowledge attention network to recognize four tachyarrhythmias, highlighting the need for deep learning approaches to consider inter-channel shared information and time-sequence dependent signals. In another study [12], a lightweight custom CNN and a transfer learning model with the MobileNet-V2 architecture were applied to segmented ECG image data that was balanced by an auxiliary classifier generative adversarial network. Chen et al. [13] developed an ensemble multiscale CNN with two modules: a Long Short-Term Memory (LSTM) network that learns from 1-D signals and a module that learns from time-frequency spectrograms from denoised ECG signals using VGGNet. The Continuous Wavelet Transform (CWT) scalogram from a single-lead ECG signal was employed in a study by Xie et al. [14]. Sarankumar et al. [15] proposed a bidirectional Gated Recurrent Unit (GRU) with an autoencoder to detect ECG arrhythmias. A two-stage ECG classifier is proposed in [16] that incorporates both ML models and CNNs. Zabihi et al. [17] proposed an innovative hybrid machine learning and deep learning approach to classify ECG beats. In a different study, spectrograms of 1D-ECG signals were created using the superlet-transform-based (SLT) method and classified with ResNet-18, as an alternative to the CWT [18]. In another study, an innovative deep learning approach was proposed that incorporates a spectral correlation function to detect atrial fibrillation [19]. Gupta et al. proposed CardioNet, combining Short-Time Fourier Transform (STFT), Gaussian window-based S-transform (ST), and Smoothed Pseudo-Wigner–Ville Distribution (SPWVD), followed by a CNN for prediction [20]. Islam et al. [21] proposed an attention, convolution, and transformer-based network to effectively capture heartbeat morphological characteristics using peak annotation on ECG signals, incorporating preprocessing algorithms such as Synthetic Minority Oversampling Technique-Edited Nearest Neighbor (SMOTE-ENN), SMOTE-Tomek, and Adaptive Synthetic (ADASYN) sampling. A hybrid deep learning framework was proposed in [22], which incorporates a specialized attention mechanism. In another approach, a novel inter-patient heartbeat classification model, G2-ResNeXt, was proposed to enhance the recognition of patients’ heart conditions by incorporating a two-fold grouping convolution (G2) into the original ResNeXt structure [23]. To tackle class imbalances in ECG data, a context-aware loss was introduced in [24], which dynamically adjusts weights according to class representation.
@@
Explainable Artificial Intelligence (XAI) methods are relatively new in the field of deep learning-based biomedical signal analysis, with only a few studies exploring various XAI techniques for ECG classification. Jo et al. [25] developed a CNN architecture with submodules to detect the presence of P waves and RR irregularities, using Gradient-weighted Class Activation Mapping (Grad-CAM) to interpret the results. Bender et al. [26] employed multiple Layer-wise Relevance Propagation (LRP) rules and Integrated Gradients (IG) for interpreting different arrhythmias. In contrast, Singh & Sharma [27] performed a more comprehensive comparison of four XAI methods, including Grad-CAM and Shapley Additive Explanations (SHAP). Goettling et al. [28] compared 13 different XAI methods, and among those techniques, the Deep Taylor Decomposition (DTD) explanation was identified as the most trustworthy. H. Le et al. [29] used modified Grad-CAM for lead-wise explanation. They employed Spearman’s rank correlation to assess the quality of the XAI approach. In another work, Ayano et al. [30] employed two model interpretability techniques, namely: Grad-CAM++ and SHAP.
Most deep learning studies, whether based on CNNs, Recurrent Neural Networks (RNNs), or a hybrid of both, usually train models using either raw ECG signals or hand-crafted features. However, some arrhythmias have identifiable features that appear as specific patterns in the signal shape; therefore, accurate interpretation requires analyzing both the signal and its morphology. A few works have merged quantitative features with deep networks and reported performance gains, but this dependence on quantitative feature inputs complicates real-time deployment on resource-constrained edge devices such as smartphones. At the same time, XAI tools in the literature are used mainly as post-hoc demonstrations of a network’s trustworthiness; they seldom feed back into the learning process itself to foster more reliable predictions. In this paper, these limitations are addressed by introducing an innovative architecture, EXGnet, which provides a reliable, trustworthy, and interpretable deep learning framework for ECG analysis. The major contributions of this paper are as follows:
  1. 1.
@@
    An innovative deep learning architecture is proposed that efficiently extracts multiresolution features from the ECG signal while minimizing computational complexity. This approach enables the model to learn both short-term and long-term features within a single network.
  2. 2.
@@
    XAI guidance during training is introduced through a Normalized Cross-Correlation (NCC)-based loss function to guide Class Activation Map (CAM) values. This strategic approach ensures that the proposed deep learning model focuses on the specific regions of ECG signals that cardiologists typically examine during their decision-making.
  3. 3.
    The ground truth CAMs for XAI guidance are automatically generated using an innovative heart rate variability-based method without the need for manual annotation.
  4. 4.
@@
    The ECG quantitative features are strategically integrated solely during the training phase, allowing the network to incorporate them for enhanced learning. Remarkably, during testing, these features are not required, yet the model demonstrates improved performance compared to when they were excluded during training.
  5. 5.
    The innovative quantitative approach for interpretability analysis, combined with qualitative visualizations, highlights the reliability and trustworthiness of the proposed network.
By bringing together these innovative components in a single framework, we offer a strong and flexible solution that works well in both real-world and remote settings. Figure 1 shows where EXGnet can be implemented in a portable health monitoring system. We evaluated our model on two publicly available and widely-used ECG arrhythmia datasets, Chapman and Ningbo, and the results show that it performs better than recent state-of-the-art methods by a clear margin. The inclusion of XAI guidance during training further improves the trustworthiness of the network by making its decisions more understandable and reliable. The rest of this paper is arranged as follows. Section 2 describes the materials and methods, including information about the datasets, how we created the ground truth CAMs, the network design, and the loss functions. Section 3 presents the experimental setup and comparative results. Section 4 includes an ablation study, an analysis of the model’s reliability, confusing cases identified in the Ningbo dataset, and discusses possible limitations. Lastly, Section 5 summarizes our findings.
@@ 2.1.1 Datasets
To evaluate the performance of EXGnet, all experiments are conducted using the publicly available database from the PhysioNet/Computing in Cardiology Challenge 2021 [31]. This database includes ECG recordings from various challenges and hospitals, totaling 45,512 recordings from the Chapman-Shaoxing [32] (Chapman dataset) and Ningbo [33] datasets, each sampled at 500 Hz for 10 seconds. The rationale behind selecting these two datasets is their widespread use in the existing literature.
The following heart conditions are focused on in this research: Sinus Rhythm (SR), Sinus Arrhythmia (SA), Sinus Tachycardia (ST), Sinus Bradycardia (SB), Atrial Fibrillation (AFib), Atrial Flutter (AF), Atrial Tachycardia (AT), and Supraventricular Tachycardia (SVT). These were selected for effective identification using the lead II signals. To enable multiclass classification, it is ensured that each ECG recording has a unique class, although some signals may belong to multiple categories. This resulted in 5,086 recordings from the Chapman dataset and 16,411 from the Ningbo dataset. Due to insufficient data, AF with AFib and AT with SVT are combined. In addition, the existing literature shows that the AF and AFib signals are often quite similar in these datasets [34]. The class distribution is detailed in Table 1. Due to the limited availability of only 15 samples for the combined AT and SVT class in the Ningbo dataset, this class was excluded from the analysis.
@@ 2.1.2 Quantitative features extraction
In our work, we utilized 17 heart rate variability based quantitative features extracted from ECG signals using the NeuroKit 0.2 Python library [35]. This library offers a comprehensive set of functions that can be applied to ECG data to extract essential features, aiding in detailed signal analysis. The following 17 features have been extracted from the ECG signal: beats per minute (BPM), mean of normal-to-normal intervals (Mean NN), standard deviation of successive differences (SDSD), standard deviation of NN intervals (SDNN), and root mean square of successive differences (RMSSD). Additional features include the coefficient of variation of successive differences (CVSD), coefficient of variation of NN intervals (CVNN), median NN intervals (Median NN), median absolute deviation of NN intervals (MAD NN), and the median coefficient of variation of NN intervals (MCV NN). We also extracted the interquartile range of NN intervals (IQR NN), the standard deviation of the RMSSD (SDRMSSD), 20th and 80th percentiles of NN intervals (Prc20 NN, Prc80 NN), minimum NN intervals (Min NN), maximum NN intervals (Max NN), and the heart rate turbulence index (HTI).
@@ 2.1.3 Data preprocessing
Figure 3: Detailed block diagram of the proposed EXGnet.